OpenStudy (anonymous):

Solve. Express your answer in simplest form. 3/4÷5 3/8 A. 6/43 B. 11/12 C. 4 1/32 D. 7 1/6

OpenStudy (anonymous):

This should help :) http://www.calculatorsoup.com/calculators/math/mixednumbers.php

OpenStudy (amorfide):

\[5\frac{ 1 }{ 4 }= \frac{ 5 }{ 1 }+\frac{ 1 }{ 4 }\] \[\frac{ a }{ b} + \frac{ c }{ d }=\frac{ (a \times d) + (c \times b) }{ b \times d }\] \[\frac{ a }{ b } \div \frac{ c }{ d } = \frac{ a }{ b } \times \frac{ d }{ c }\] \[\frac{ a }{ b} \times \frac{ c }{ d} = \frac{ a \times d }{ b \times c}\]

OpenStudy (amorfide):

use this sort of logic for your question

OpenStudy (amorfide):

then to simplify, let's say we are simplifying \[\frac{ 4 }{ 6 }\] you look for the factors of each number, 4 and 6 so the factors of 4 1,2,4 factors of 6 1,2,3,6 you look for the highest factor they have in common, which is 2 so you divide 4 by 2 and divide 6 by 2 to get \[\frac{ 2 }{ 3 }\]